OpenStudy (anonymous):

Solve 25^x = 5^(x^2-3) for x

OpenStudy (anonymous):

\[25^{x}=5^{x^{2}-3}\] Solve for x

jimthompson5910 (jim_thompson5910):

25 = 5^2

jimthompson5910 (jim_thompson5910):

so we can say \[\large 25^{x}=5^{x^{2}-3}\] \[\large (5^2)^{x}=5^{x^{2}-3}\] \[\large 5^{2x}=5^{x^{2}-3}\] and because the bases are equal, the exponents are equal, meaning that \[\large 2x = x^2 - 3\]

jimthompson5910 (jim_thompson5910):

do you see what to do from here?

OpenStudy (anonymous):

Yes thank you.

jimthompson5910 (jim_thompson5910):

ok great, yw